Our engineering team continued to monitor our storage platform for the rest of the day, and performance returned to normal.
At this time we consider everything to be operating optimally, and as such we are closing this case.
Our engineering team worked with our upstream storage platform vendor to improve performance, and this has been greatly improved to close to normal levels over the past two hours.
Background maintenance operations continue, and we will be working further on this outside of core UK/European business hours this evening.
We are seeing degraded performance on our storage cluster which is impacting a number of services. A rebalancing operation is in place and we hope to restore normal service soon.